For what values of m is the line y=mx, a tangent to the parabola y= x^2 - 8x + 25? Thank you :)
Answer by rothauserc(4718)

y = f(x) = x^2 -8x +25
:
let (a,f(a)) be a point on f(x)
:
the tangent line at (a,f(a)) is
:
y = f'(a)(x-a)+f(a)
:
f'(x) = 2x -8, f(a) = a^2 -8a +25
:
f'(a) = 2a -8
:
y = 2a -8(x-a) +a^2 -8a +25
:
y = -a^2 +2ax -8x +25
:
we want y = mx, then
:
-a^2 +25 = 0
:
a^2 = 25
:
a = 5 or -5
:
m = f'(a) = 2a -8
:
m = 2 or -18
: